could someone please show me how to find the vertex, line of symmetry, the maximum or minimum value of the quadratic function and graph the function. f(x)=5-x^2
Is the value, f(0)=5 a minimum or maximum?

Answer by swincher4391(1107)

The vertex can be found by using the f(x) = ax^2 + bx + c where the vertex is (-b/2a, f(-b/2a) So let's find -b/2a.
-b/2a= -0/2(-1) = 0
And f(0) = 5.
Thus the vertex is at (0,5).
In a parabola, the line of symmetry is at the x-value of the vertex. So the line of symmetry is the y-axis or x=0.
To graph this, take the parent function y=x^2, a parabola, flip it upside down, and shift it up 5 spaces. Notice that your vertex is indeed at (0,5) at the top of the parabola. Thus f(0) = 5 is a maximum.
Hope this helped!
Please e-mail me at onlyyourmathtutor@yahoo.com if you have any other questions.